# Potential-energy gating for robust state estimation in bistable stochastic systems

### v1

@@ III.1 Motivation: energy-dependent observation reliability
where $g\geq 0$ is the gating sensitivity. When $g=0$, the standard filter is recovered. As $g$ increases, the filter progressively discounts observations when the state estimate is in a high-energy region.
@@ IV.3 Goldilocks zone for hyperparameters
Figure 3: Sensitivity of PG-EKF performance to hyperparameters $\lambda$ and $g$ across four potential topologies ($\alpha=0.6,1.0,1.41,2.0$). A broad “Goldilocks zone” ($\lambda\in[0.05,0.5]$, $g\in[5,30]$, shaded region) provides near-optimal RMSE reduction regardless of the specific potential shape. The robustness of this plateau means that approximate parameter knowledge is sufficient for effective gating.
@@ V.4 Filter performance on NGRIP data
Table 3: PG-EKF improvement (%) on NGRIP data ($\hat{\gamma}=-0.109$, 2020 MC replications). Values are mean $\pm$ SE. At $0\%$ outlier, $\mathrm{SE}=0$ because the same clean trajectory produces identical filter output.
@@ VII Conclusion
The method is applicable to any system where the existence and approximate location of metastable states can be specified a priori from domain knowledge, experimental evidence, or preliminary data analysis. Natural extensions include multivariate state spaces, online estimation of potential parameters, and coupling with data-driven methods for systems where the potential form is not known analytically.

### v2

@@ I Introduction
The need for physics-based filtering is especially acute in domains where only a single realization of the process is available—paleoclimate records, seismological catalogs, or financial time series of rare regime changes. In such non-ergodic or data-scarce settings, purely statistical methods cannot reliably learn the noise structure from the data alone, because the number of observed transitions is too small for robust parameter estimation. Incorporating prior knowledge about the energy landscape offers a principled alternative: the physics constrains the filter where the data cannot.
@@ III.1 Motivation: energy-dependent observation reliability
where $g\geq 0$ is the gating sensitivity. When $g=0$, the standard filter is recovered. As $g$ increases, the filter progressively discounts observations when the state estimate is in a high-energy region. The linear form 1+gV(x)1+g\,V(x) is the simplest choice that satisfies two requirements: (i) $R(x)\to R_{0}$ at the well minima where $V(x)=0$, recovering nominal trust, and (ii) $R(x)$ grows monotonically with potential energy, reflecting decreasing measurement reliability near the barrier. Formally, it is the first-order Taylor expansion of any monotonically increasing gating function $\phi(V)$ satisfying $\phi(0)=1$, with $g=\phi^{\prime}(0)$. Higher-order forms (e.g., exponential gating $R_{0}\,e^{gV}$) are possible but offer no qualitative advantage: the misspecification analysis in Sec. IV.4 shows that performance is dominated by the topological structure of the potential rather than the precise functional form of the gating.
@@ IV.3 Goldilocks zone for hyperparameters
Figure 3: Sensitivity of PG-EKF performance to the gating sensitivity $g$ across four potential topologies with increasing barrier height: $\alpha=0.6$ ($\Delta V=0.09$), $\alpha=1.0$ ($\Delta V=0.25$), $\alpha=1.41$ ($\Delta V=0.50$), $\alpha=2.0$ ($\Delta V=1.00$), where $\Delta V=\alpha^{2}/4\beta$ is the barrier height. A: RMSE vs. $g$ for each $\alpha$; a broad “Goldilocks zone” ($g\in[5,50]$, shaded) provides near-optimal performance regardless of potential shape. B: Optimal $g$ as a function of barrier height, confirming that a single default ($g=20$) works across topologies. C: Percentage RMSE degradation from each topology’s optimum, showing that the Goldilocks zone corresponds to $5\%$ degradation across all barrier heights. The regularization strength is fixed at $\lambda=0.5$ throughout.
@@ V.4 Filter performance on NGRIP data
Table 3: PG-EKF RMSE improvement (%) over the standard EKF on NGRIP $\delta^{18}$O data, as a function of sub-sampled record length NN (rows) and synthetic outlier contamination fraction (columns). The asymmetric Ginzburg-Landau model uses $\hat{\gamma}=-0.109$. Each cell reports the mean improvement across 2020 Monte Carlo replications; the $\pm$ values denote the standard error of the mean ($\mathrm{SE}=\sigma/\sqrt{20}$, where $\sigma$ is the standard deviation across replications). At $0\%$ outlier contamination, $\mathrm{SE}=0$ because the deterministic EKF applied to the same clean trajectory produces identical output in every replication. Improvement increases monotonically with outlier fraction at fixed $N$, and is approximately stable across NN at fixed outlier fraction (ANOVA: Sec. V.4).
@@ VII Conclusion
The method is applicable to any system where the existence and approximate location of metastable states can be specified a priori from domain knowledge, experimental evidence, or preliminary data analysis. It is especially valuable in non-ergodic or data-scarce settings—such as paleoclimate records, seismological catalogs, or financial time series of rare regime changes—where only a single realization is available and the number of observed transitions is too small for purely statistical approaches to learn the noise structure reliably. Natural extensions include multivariate state spaces, online estimation of potential parameters, and coupling with data-driven methods for systems where the potential form is not known analytically.